# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

### v1

@@ A.2 First order results
$\mathcal{J}^{(1)}=\sqrt{\bar{n}}\left[\begin{pmatrix}0\\ \mathcal{J}_{+1,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}\mathcal{J}_{-1,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\right]\,,$ (52)
@@
$\mathcal{J}_{+1,\downarrow}^{(1)}(s_{p})=\mathcal{J}_{-1,\uparrow}^{(1)}(-s_{p})=\sqrt{\frac{1+s_{p}}{2}}\frac{\hbar\Omega_{R}}{2}\,.$ (53)
Using this expression and following the procedure outlined in Sec. A.1, one immediately finds that $\mu^{(1)}=\Delta\mu^{(1)}=0$. Furthermore, the first order correction to the order parameter, satisfying both Eq. (45) and the orthogonality conditions (51), takes the form
$\begin{split}\Psi_{0}^{(1)}={}&{}\sqrt{\bar{n}}\bigg{[}\begin{pmatrix}\tilde{\Psi}_{+3,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{+1,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{-1,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-3,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}\bigg{]}\,,\end{split}$ (54)
where the coefficients are given by
$\tilde{\Psi}_{+1,\downarrow}^{(1)}(s_{p})$ ${}=\tilde{\Psi}_{-1,\uparrow}^{(1)}(-s_{p})=\sqrt{\frac{1+s_{p}}{2}}\frac{N_{1}^{(1)}}{2D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,,$ (55a)
$\begin{split}\tilde{\Psi}_{+3,\uparrow}^{(1)}(s_{p})&{}=\tilde{\Psi}_{-3,\downarrow}^{(1)}(-s_{p})\\ &{}=\sqrt{\frac{1-s_{p}}{2}}\frac{(1+s_{p})N_{3}^{(1)}}{2D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,.\end{split}$ (55b)
The explicit forms of $D_{1}$, $D_{2}$, $N_{1}^{(1)}$, and $N_{3}^{(1)}$ are given in Appendix B. These results are obtained by inserting the Ansatz (54) into Eq. (45), grouping terms with the same oscillatory behavior, and equating coefficients on both sides of the equation. Once $\Psi_{0}^{(1)}$ is known, any observable can be computed to first order in $\hbar\Omega_{R}/4E_{R}$. Among those discussed in Sec. IV.3, only the density profiles [see Eqs. (25), (26), and (27)] and the associated contrasts exhibit nonzero first order corrections.
@@ A.3 Second order results
$\begin{split}\mathcal{J}^{(2)}={}&{}\sqrt{\bar{n}}\bigg{[}\begin{pmatrix}\mathcal{J}_{+5,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{5\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{+3,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\mathcal{J}_{+1,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{-1,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\mathcal{J}_{-3,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{-5,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-5\mathrm{i}X}\bigg{]}\,.\end{split}$ (56)
@@
The coefficients in this expression can be derived from the general formula (47), using the results of the previous section. They satisfy the symmetry relations $\mathcal{J}_{+5,\uparrow}^{(2)}(s_{p})=\mathcal{J}_{-5,\downarrow}^{(2)}(-s_{p})$, $\mathcal{J}_{+3,\downarrow}^{(2)}(s_{p})=\mathcal{J}_{-3,\uparrow}^{(2)}(-s_{p})$, and $\mathcal{J}_{+1,\uparrow}^{(2)}(s_{p})=\mathcal{J}_{-1,\downarrow}^{(2)}(-s_{p})$. Applying the method of Sec. A.1, one obtains the second order corrections to the chemical potentials $\mu^{(2)}$ and $\Delta\mu^{(2)}$, which are reported in the main text [see second term on the right hand side of Eq. (33) and Eq. (34), respectively]. The second order correction to the order parameter, derived in the same manner as the first order one (see Sec. A.2), reads
$\begin{split}\Psi_{0}^{(2)}={}&{}\sqrt{\bar{n}}\bigg{[}\begin{pmatrix}\tilde{\Psi}_{+5,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{5\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{+3,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-1,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{-3,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-5,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-5\mathrm{i}X}\bigg{]}\,,\end{split}$ (57)
with the coefficients given by
@@
$\begin{split}\tilde{\Psi}_{+1,\uparrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-1,\downarrow}^{(2)}(-s_{p})\\ &{}=\sqrt{\frac{1+s_{p}}{2}}\frac{N_{1}^{(2)}}{4D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,\end{split}$ (58a)
$\begin{split}\tilde{\Psi}_{+3,\downarrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-3,\uparrow}^{(2)}(-s_{p})\\ &{}=-\sqrt{\frac{1-s_{p}}{2}}\frac{(1+s_{p})N_{3}^{(2)}}{4D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,\end{split}$ (58b)
$\begin{split}\tilde{\Psi}_{+5,\uparrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-5,\downarrow}^{(2)}(-s_{p})\\ &{}=\sqrt{\frac{1+s_{p}}{2}}\frac{(1-s_{p}^{2})N_{5}^{(2)}}{4D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,.\end{split}$ (58c)
@@
The quantities $D_{1}$, $D_{2}$, $N_{1}^{(2)}$, $N_{3}^{(2)}$, and $N_{5}^{(2)}$ are given in Appendix B. The structure of Eqs. (58) respects the orthogonality constraints (51). With the expression for $\Psi_{0}^{(2)}$ at hand, the second order corrections to observables can be readily computed. Selected results are presented in Sec. IV.3 of the main text. Finally, it is worth noting that all the perturbative expressions reported in this appendix reduce to those of Ref. [58] when $s_{p}=0$.
@@ Appendix B Coefficients of perturbative formulas
In this Appendix we give the explicit forms of the coefficients introduced in Secs. IV.3, A.2, and A.3. These coefficients appear in various perturbative formulas.
@@
×${}\times\left[\left(4E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
@@
×${}\times\left[\left(16E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
@@
The coefficients entering the numerators of the observables computed in Sec. IV.3 are
@@
$N_{k_{1}}={}$ ${}4E_{R}^{2}\left(2E_{R}+G_{ss}\right)^{2}\left(4E_{R}^{2}+2E_{R}G_{dd}+G_{dd}^{2}\right)$
${}+4E_{R}\big{[}4E_{R}^{3}\left(4E_{R}+3G_{ss}\right)G_{ss}+\big{(}8E_{R}^{4}$
${}+16E_{R}^{3}G_{ss}+4E_{R}^{2}G_{ss}^{2}-3E_{R}G_{ss}^{3}-G_{ss}^{4}\big{)}G_{dd}\big{]}s_{p}^{2}$
${}-G_{ss}^{2}\big{[}\left(16E_{R}^{3}-16E_{R}^{2}G_{ss}+2E_{R}G_{ss}^{2}+G_{ss}^{3}\right)G_{ss}$
${}-4E_{R}^{2}\left(2E_{R}+G_{ss}\right)G_{dd}\big{]}s_{p}^{4}-2E_{R}G_{ss}^{5}s_{p}^{6}\,,$ (63)
$N_{\langle\sigma_{z}\rangle}={}$ ${}4E_{R}^{2}\big{[}2E_{R}\left(4E_{R}^{2}+6G_{ss}E_{R}+G_{ss}^{2}\right)$
@@
$N_{E}={}$ ${}2E_{R}\left(2E_{R}+G_{ss}\right)\left(4E_{R}+G_{dd}\right)$
@@
$N_{\mu}={}$ ${}-4E_{R}^{3}\left(2E_{R}+G_{ss}\right)^{2}\left(8E_{R}^{2}+4E_{R}G_{dd}+G_{dd}^{2}\right)$
@@
$N_{v_{x}}={}$ ${}4E_{R}^{2}(2E_{R}+G_{ss})\big{[}\left(4E_{R}^{2}+6E_{R}G_{ss}+G_{ss}^{2}\right)G_{ss}$
${}+\left(4E_{R}^{2}+8E_{R}G_{ss}+3G_{ss}^{2}\right)G_{dd}$
${}+\left(2E_{R}+G_{ss}\right)G_{dd}^{2}\big{]}-2E_{R}G_{ss}\big{[}4\big{(}4E_{R}^{3}$
${}+5E_{R}^{2}G_{ss}+3E_{R}G_{ss}^{2}+G_{ss}^{3}\big{)}G_{ss}+\big{(}8E_{R}^{3}$
${}+16E_{R}^{2}G_{ss}+8E_{R}G_{ss}^{2}+G_{ss}^{3}\big{)}G_{dd}\big{]}s_{p}^{2}$
${}+2E_{R}G_{ss}^{3}[4\left(E_{R}+G_{ss}\right)G_{ss}$
${}-\left(2E_{R}+G_{ss}\right)G_{dd}]s_{p}^{4}+G_{ss}^{6}s_{p}^{6}\,.$ (68)
@@
×${}\times\left[\left(4E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}$

### v2

@@ A.2 First order results
$\mathcal{J}^{(1)}=\begin{pmatrix}0\\ \mathcal{J}_{+1,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}\mathcal{J}_{-1,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\,$, (66)
@@
$\mathcal{J}_{+1,\downarrow}^{(1)}(s_{p})=\mathcal{J}_{-1,\uparrow}^{(1)}(-s_{p})=\sqrt{\frac{\bar{n}(1+s_{p})}{2}}\frac{\hbar\Omega_{R}}{2}\,.$ (67)
In addition, following the procedure outlined in Sec. A.1, one immediately finds that $\bar{\varepsilon}^{(1)}=0$, meaning that $\mu_{d}^{(1)}=\mu_{s}^{(1)}=0$ and $\alpha_{c}^{(1)}=\alpha_{1}^{(1)}=0$. The first order correction to the order parameter, satisfying both Eq. (59) and the orthogonality conditions (65), takes the form
$\begin{split}\Psi_{0}^{(1)}={}&{}\begin{pmatrix}\tilde{\Psi}_{+3,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{+1,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{-1,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-3,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}\,.\end{split}$ (68)
To determine the coefficients in the above expression, we insert the Ansatz (68) into Eq. (59), group terms with the same oscillatory behavior, and equate coefficients on both sides of the equation. The final results are functions of $s_{p}$, $\alpha_{c}$, and $\alpha_{1}$. Here we provide the explicit expressions when $\alpha_{c}$ and $\alpha_{1}$ are taken equal to the optimal values given above:
$\tilde{\Psi}_{+1,\downarrow}^{(1)}(s_{p})$ ${}=\tilde{\Psi}_{-1,\uparrow}^{(1)}(-s_{p})=\sqrt{\frac{\bar{n}(1+s_{p})}{2}}\frac{N_{1}^{(1)}}{2D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,$, (69a)
$\begin{split}\tilde{\Psi}_{+3,\uparrow}^{(1)}(s_{p})&{}=\tilde{\Psi}_{-3,\downarrow}^{(1)}(-s_{p})\\ &{}=\sqrt{\frac{\bar{n}(1-s_{p})}{2}}\frac{(1+s_{p})N_{3}^{(1)}}{2D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,.\end{split}$ (69b)
The explicit forms of $D_{1}$, $D_{2}$, $N_{1}^{(1)}$, and $N_{3}^{(1)}$ are given in Appendix B. Once $\Psi_{0}^{(1)}$ is known, any observable can be computed to first order in $\hbar\Omega_{R}/4E_{R}$. Among those discussed in Sec. IV.2, only the density profiles [see Eqs. (33), (34), (35a), and (35c)] and the associated contrasts exhibit nonzero first order corrections.
@@ A.3 Second order results
$\begin{split}\mathcal{J}^{(2)}={}&{}\begin{pmatrix}\mathcal{J}_{+5,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{5\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{+3,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\mathcal{J}_{+1,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{-1,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\mathcal{J}_{-3,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{-5,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-5\mathrm{i}X}\,.\end{split}$ (70)
@@
The coefficients in this expression can be derived from the general formula (61), using the results for the first order corrections to the various quantities. Applying the method of Sec. A.1, one obtains the second order corrections to the average energy density $\bar{\varepsilon}^{(2)}$, the mean chemical potentials $\mu_{d}^{(2)}$, the chemical potential semi-difference $\mu_{s}^{(2)}$, and the dimensionless wave vectors $\alpha_{c}^{(2)}$ and $\alpha_{1}^{(2)}$. These are all reported in the main text [see second term on the right hand side of Eqs. (41), (42), (43), (38), and (39)]. The second order correction to the order parameter, derived in the same manner as the first order one (see Sec. A.2), reads
$\begin{split}\Psi_{0}^{(2)}={}&{}\begin{pmatrix}\tilde{\Psi}_{+5,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{5\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{+3,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-1,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{-3,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-5,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-5\mathrm{i}X}\,.\end{split}$ (71)
As for the first order corrections, we give the explicit expressions of the coefficients in Eq. (71) only for $\alpha_{c}$ and $\alpha_{1}$ equal to their optimal values:
[…]
@@
$\begin{split}\tilde{\Psi}_{+5,\uparrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-5,\downarrow}^{(2)}(-s_{p})\\ &{}=\sqrt{\frac{\bar{n}(1+s_{p})}{2}}\frac{(1-s_{p}^{2})N_{5}^{(2)}}{4D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,.\end{split}$ (72c)
@@
The quantities $D_{1}$, $D_{2}$, $N_{1}^{(2)}$, $N_{3}^{(2)}$, and $N_{5}^{(2)}$ are given in Appendix B. The structure of Eqs. (72) respects the orthogonality constraints (65). With the expression for $\Psi_{0}^{(2)}$ at hand, the second order corrections to observables can be readily computed. In addition to those mentioned earlier, in Sec. IV.2 of the main text we present the results for the total and bare spin density profiles [see Eqs. (33), (34), (35b), and (35d)], the average bare spin polarization density $\bar{s}_{z}$ [Eq. (40)], and the fringe velocity $v$ [Eq. (44)]. Finally, it is worth noting that all the perturbative expressions reported in this appendix reduce to those of Ref. [60] when $s_{p}=0$.
@@ Appendix B Coefficients of perturbative formulas
In this Appendix we give the explicit forms of the coefficients introduced in Secs. IV.2, A.2, and A.3. These coefficients appear in various perturbative formulas.
@@
${}\times\left[\left(4E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
@@
${}\times\left[\left(16E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
@@
The coefficients entering the numerators of the observables computed in Sec. IV.2 are
@@
$N_{k_{c}}={}$ ${}4E_{R}^{2}(2E_{R}+G_{ss})\big[\left(4E_{R}+G_{ss}\right)G_{ss}$
${}+2\left(E_{R}+G_{ss}\right)G_{dd}\big]$
${}-2E_{R}G_{ss}^{2}\big[\left(4E_{R}+3G_{ss}\right)G_{ss}$
${}-\left(2E_{R}+G_{ss}\right)G_{dd}\big]s_{p}^{2}-G_{ss}^{5}s_{p}^{4}\,$, (77)
$N_{k_{1}}={}$ ${}4E_{R}^{2}(2E_{R}+G_{ss})^{2}\left(4E_{R}^{2}+2E_{R}G_{dd}+G_{dd}^{2}\right)$
${}-4E_{R}G_{ss}\big[2E_{R}^{2}G_{ss}^{2}-\big(4E_{R}^{3}-3E_{R}G_{ss}^{2}$
${}-G_{ss}^{3}\big)G_{dd}\big]s_{p}^{2}-G_{ss}^{4}\left(4E_{R}^{2}-2E_{R}G_{ss}-G_{ss}^{2}\right)s_{p}^{4}\,$, (78)
$N_{\bar{s}_{z}}={}$ ${}4E_{R}^{2}\big[2E_{R}\left(4E_{R}^{2}+6G_{ss}E_{R}+G_{ss}^{2}\right)$
@@
$N_{\bar{\varepsilon}}={}$ ${}2E_{R}\left(2E_{R}+G_{ss}\right)\left(4E_{R}+G_{dd}\right)$
@@
$N_{\mu_{d}}={}$ ${}-4E_{R}^{3}\left(2E_{R}+G_{ss}\right)^{2}\left(8E_{R}^{2}+4E_{R}G_{dd}+G_{dd}^{2}\right)$
@@
$N_{v}={}$ ${}2E_{R}G_{ss}\left(4E_{R}^{2}+6E_{R}G_{ss}+G_{ss}^{2}\right)(2E_{R}+G_{ss})$
${}+2E_{R}(2E_{R}+3G_{ss})(2E_{R}+G_{ss})^{2}G_{dd}$
${}+2E_{R}(2E_{R}+G_{ss})^{2}G_{dd}^{2}$
${}-\Big[2G_{ss}^{3}\left(4E_{R}^{2}+5E_{R}G_{ss}+2G_{ss}^{2}\right)$
${}+(2E_{R}+G_{ss})^{2}G_{ss}^{2}G_{dd}\Big]s_{p}^{2}+G_{ss}^{5}s_{p}^{4}\,.$ (83)
@@
${}\times\left[\left(4E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}$